# Degree two Gopakumar–Vafa invariants of local curves

### v2

@@ 3.3. Fibrewise GV invariants
$P(T_{2,n})=-a(q-q^{-1})P(T_{2,n-2})+a^{2}P(T_{2,n-2})\text{ for }n\geq 2.$
This last equation is the standard recursion relation for HOMFLY polynomials of torus knots, and so $P(T_{2,n})$ is the HOMFLY polynomial of the torus knot $T_{2,n}$. We will use the specialization $(q/a)^{n-1}P(T_{2,n})|_{a=0}$ for $2\leq n\leq 5$:
@@
Let $C_{a}\to C$ be the spectral cover corresponding to a point $a\in S_{\lambda}\cap\mathbb{P}\operatorname{H}^{0}(L^{\otimes 2})$. For any partition $\lambda$, the spectral curve $C_{a}$ is reduced. Hence, by [MS13, MY14], the fibrewise contribution $n_{g}(\lambda)$ is computed by the Euler characteristics of the Hilbert schemes of points on $C_{a}$. The latter is then computed by using certain HOMFLY polynomials. More precisely, [MS13, MY14] gives the first, and [Mau16, OS12] gives the second equality in the following sequence of equalities:
@@
$=(1-q^{2})^{-\chi(C)}\prod_{i=1}^{r(\lambda)}[(q/a)^{i-1}P(T_{2,\lambda_{i}})]_{a=0}$
$=q^{-\chi(C)}(q^{-1}-q)^{-\chi(C)}\prod_{i=1}^{r(\lambda)}[(q/a)^{i-1}P(T_{2,\lambda_{i}})]_{a=0}.$
where we denote by $r(\lambda)$ the number of rows of $\lambda$. We give a little more detail, in order to explain the sign $(-1)^{g}$ appearing in the LHS of (3.4). Note that $102g(C_{a})-2=$ 10 for all $a$. By [MS13, MY14], we have the following (cf. [MT18][Equation (4.4)]):
@@ 3.4. Contributions from the nearby hyperplane
$\quad-2\cdot 542-3\cdot(128)-2\cdot 968-1\cdot(-1012)$
@@ 3.5. GV invariants for twisted Higgs bundles
Let $\hat{M}$ denote the moduli space of $L$-twisted $\operatorname{PGL}_{n}$-Higgs bundles. By [dCHM12][Theorems 1.2.10 and 4.2.2], we can see that
@@
$\operatorname{H}_{\leq 2}^{*}(\hat{M})=\operatorname{H}_{\leq 3}^{*}(\hat{M})=\langle\alpha,\beta,\psi_{i}\rangle_{1\leq i\leq 4},$
$\operatorname{H}_{\leq 4}^{*}(\hat{M})=\operatorname{H}_{\leq 5}^{*}(\hat{M})=\langle\operatorname{H}_{\leq 2}^{*}(\hat{M}),\alpha^{2},\alpha\beta,\alpha\psi_{i},\psi_{i}\psi_{j}\rangle_{1\leq i,j,\leq 4},$
$\operatorname{H}_{\leq 6}^{*}(\hat{M})=\operatorname{H}_{\leq 7}^{*}(\hat{M})=\langle\operatorname{H}_{\leq 2}^{*}(\hat{M}),\alpha^{3},\alpha^{2}\beta,\alpha^{2}\psi_{i}\rangle_{1\leq i\leq 4},$
$\operatorname{H}_{\leq 8}^{*}(\hat{M})=\langle\operatorname{H}_{\leq 6}^{*}(\hat{M}),\alpha^{4}\rangle,$
where the cohomological degrees of $\alpha,\beta,\psi_{i}$ are 2, 4, 3, respectively. Here we have followed the notation and normalisation conventions of [dCHM12], regarding the perverse filtration on $\operatorname{H}^{*}(\hat{M})$. Precisely:
@@
where $a=\dim(\tilde{B}_{L})=4g(C)-1$. Combined with the isomorphism $\operatorname{H}^{*}(M)\cong\operatorname{H}^{*}(\hat{M})\otimes\operatorname{H}^{*}(\operatorname{Pic}^{0}(C))$, we obtain
$\sum_{i}\chi({}^{p}\mathcal{H}^{i}(^{*}{\mathbf{R}h_{*}\operatorname{IC}_{M}}))q^{i}=-(q^{-4}-2q^{-2}+4-2q^{2}+q^{4})(q^{-1/2}+q^{1/2})^{4},$
from which we can easily read off the integers $n_{g,2[C]}(\operatorname{Tot}_{C}(L))$. Note that the minus sign in the right hand side comes from the fact that $\dim M=8g(C)-3$ is odd. ∎
@@ 4. Higher genus
For simplicity, we write $M=M_{L}(2,1)$ and $\hat{M}=\hat{M}_{L}(2,1)$. A smooth spectral curve has genus $24g(C)-$ 2. Hence, the degree of the Laurent polynomial $\sum_{i\in\mathbb{Z}}\chi({}^{p}\mathcal{H}^{i}(\mathbf{R}h_{*}\operatorname{IC}_{M}))q^{i}$ is $24g(C)-$ 2. The coefficient of $q^{4g(C)-2}$ is
@@
To determine the GV invarinats for small $g$, recall that we have an isomorphism
@@
From this, we can see that $n_{g,2[C]}(\operatorname{Tot}_{C}(L))=0$ for $g\leq g(C)-1$. Furthermore, we have $n_{g,2[C]}(\operatorname{Tot}_{C}(L))=-\chi(\hat{M})$. This can be computed as follows: Let $P_{t}(M)$ denote the Poincare polynomial of $M$. We can compute $P_{-t}(M)$ by substituting $u=v=-1$ in [CDP11, Equation (A.7)]:

### v3

@@ 3.3. Fibrewise GV invariants
$P(T_{2,n})=-a(q-q^{-1})P(T_{2,n-1})+a^{2}P(T_{2,n-2})\text{ for }n\geq 2.$
This last equation is the standard recursion relation for HOMFLY polynomials of torus knots, and so $P(T_{2,n})$ is the HOMFLY polynomial of the torus knot $T_{2,n}$ (see [OS12, Section 2]). We will use the specialization $(q/a)^{n-1}P(T_{2,n})|_{a=0}$ for $2\leq n\leq 5$:
@@
Let $C_{a}\to C$ be the spectral cover corresponding to a point $a\in S_{\lambda}\cap\mathbb{P}\operatorname{H}^{0}(L^{\otimes 2})$. For any partition $\lambda$, the spectral curve $C_{a}$ is reduced. Hence, by [MS13, MY14], the fibrewise contribution $n_{g}(\lambda)$ is computed by the Euler characteristics of the Hilbert schemes of points on $C_{a}$. The latter is then computed by using certain HOMFLY polynomials. More precisely, [MS13, MY14] gives the first, and [OS12, Mau16], gives the second equality in the following sequence of equalities:
@@
$=(1-q^{2})^{-\chi(C_{a})}\prod_{i=1}^{r(\lambda)}[(q/a)^{\lambda_{i}-1}P(T_{2,\lambda_{i}})]_{a=0}$
$=q^{-\chi(C_{a})}(q^{-1}-q)^{-\chi(C_{a})}\prod_{i=1}^{r(\lambda)}[(q/a)^{\lambda_{i}-1}P(T_{2,\lambda_{i}})]_{a=0}.$
where we denote by $r(\lambda)$ the number of rows of $\lambda$. We give a little more detail, firstly in order to explain the sign $(-1)^{g}$ appearing in the LHS of (3.4). Note that $102g(C_{a})-2=$ 10 for all $a$. By [MS13, MY14], we have the following (cf. [MT18][Equation (4.4)]):
@@
Let us also explain the second equality. Let $\{p_{i}\}$ be the set of singular points of $C_{a}$. By [OS12, Conjecture 1] (which is proved in [Mau16]), we have
[…]
where $\mu_{i}$ denotes the Milnor number of the singularity $p_{i}$ and $L_{i}$ denotes the link of $C_{a}$ at $p_{i}$. In our case, the singular points of $C_{a}$ are parametrised by the rows of the partition $\lambda=(\lambda_{i})$, and the type of the singular point corresponding to $\lambda_{i}$ is $\{y^{2}=x^{\lambda_{i}}\}$. As a result, we have $\mu_{i}=\lambda_{i}-1$ and $L_{i}=T_{2,\lambda_{i}}$; these equalities imply the second equality in (3.4).
@@ 3.4. Contributions from the nearby hyperplane
$\quad-2\cdot 542-3\cdot(-128)-2\cdot 968-1\cdot(-1012)$
@@ 3.5. GV invariants for twisted Higgs bundles
Let $\hat{M}$ denote the moduli space of $L$-twisted $\operatorname{PGL}_{n}$-Higgs bundles.
We first describe a basis of $\operatorname{H}^{i}(\hat{M})$ for each $i$. By [Moz12, Subsection 6.3], adopting the notation of loc. cit., the Poincare polynomial of $\hat{M}$ is
(3.5) $P_{y}(\hat{M})=P^{(1)}_{2,2}(y)/(1+y)^{4}=2y^{8}+4y^{7}+8y^{6}+4y^{5}+2y^{4}+4y^{3}+y^{2}+1.$
On the other hand, the cohomology ring $\operatorname{H}^{*}(\hat{M})$ has the following multiplicative generators [dCHM12, HT03, HT04]:
$\alpha\in\operatorname{H}^{2}(\hat{M}),\quad\beta\in\operatorname{H}^{4}(\hat{M}),\quad\psi_{i}\in\operatorname{H}^{3}(\hat{M}),\quad i=1,\cdots,4.$
Using the relations in [dCHM12, Theorem 1.2.10, Proposition 4.2.1] and the equation (3.5), we have that
$\operatorname{H}^{0}(\hat{M})=\langle 1\rangle,\quad\operatorname{H}^{1}(\hat{M})=0,\quad\operatorname{H}^{2}(\hat{M})=\langle\alpha\rangle,\quad\operatorname{H}^{3}(\hat{M})=\langle\psi_{i}\rangle_{1\leq i\leq 4},$
$\operatorname{H}^{4}(\hat{M})=\langle\alpha^{2},\beta\rangle,\quad\operatorname{H}^{5}(\hat{M})=\langle\alpha\psi_{i}\rangle_{1\leq i\leq 4},\quad\operatorname{H}^{6}(\hat{M})=\langle\psi_{i}\psi_{j},\alpha^{3},\alpha\beta\rangle_{1\leq i<j\leq 4},$
$\operatorname{H}^{7}(\hat{M})=\langle\alpha^{2}\psi_{i}\rangle_{1\leq i\leq 4},\quad\operatorname{H}^{8}(\hat{M})=\langle\alpha^{4},\alpha\gamma\rangle,$
where we put $\gamma\coloneqq-2(\psi_{1}\psi_{3}+\psi_{2}\psi_{4})$. In the following, we only explain the above equalities for $\operatorname{H}^{7}(\hat{M})$ and $\operatorname{H}^{8}(\hat{M})$. The arguments are similar and simpler for lower cohomological degrees. For $\operatorname{H}^{7}(\hat{M})$, we have
$\operatorname{H}^{7}(\hat{M})=\langle\beta\psi_{i},\alpha^{2}\psi_{i}\rangle_{1\leq i\leq 4}=\langle\alpha^{2}\psi_{i}\rangle_{1\leq i\leq 4},$
where the second equality follows from [dCHM12, Proposition 4.2.1]. On the other hand, by (3.5), we also know that $\dim\operatorname{H}^{7}(\hat{M})=4$. Hence, we can conclude that $\{\alpha^{2}\psi_{i}\}_{1\leq i\leq 4}$ is a basis of $\operatorname{H}^{7}(\hat{M})$. For $\operatorname{H}^{8}(\hat{M})$, we have
(3.6) $\operatorname{H}^{8}(\hat{M})$ $=\langle\alpha^{4},\alpha^{2}\beta,\alpha\psi_{i}\psi_{j},\beta^{2}\rangle_{1\leq i<j\leq 4}$
$=\langle\alpha^{4},\alpha^{2}\beta,\alpha\psi_{1}\psi_{2},\alpha\psi_{1}\psi_{4},\alpha\psi_{2}\psi_{3},\alpha\psi_{3}\psi_{4},\alpha(\psi_{1}\psi_{3}-\psi_{2}\psi_{4}),\alpha\gamma,\beta^{2}\rangle$
$=\langle\alpha^{4},\alpha\gamma\rangle.$
To see the third equality, we claim that
$\alpha\in I^{0}_{3},\quad\beta^{2}\in I^{2}_{1},\quad\frac{\alpha^{2}\beta}{2}+2\alpha\gamma\in I^{2}_{1},$
where $I^{a}_{b}\subset\mathbb{Q}[\alpha,\beta,\gamma]$ is an ideal defined in [dCHM12, Definition 1.2.8] for $a,b\geq 0$. Indeed, we have
$\alpha=\rho^{0}_{1,0,0},\quad\beta^{2}=\rho^{3}_{0,2,0},\quad\frac{\alpha^{2}\beta}{2}+2\alpha\gamma=\rho^{2}_{2,1,0},$
where $\rho^{c}_{r,s,t}\in I^{a}_{b}$ are elements defined in loc. cit.. Hence, we have that $\beta^{2}=0$ and $\alpha^{2}\beta\in\langle\alpha^{4},\alpha\gamma\rangle$. Furthermore, since we have
(3.7) $\psi_{1}\psi_{2},\psi_{1}\psi_{4},\psi_{2}\psi_{3},\psi_{3}\psi_{4},(\psi_{1}\psi_{3}-\psi_{2}\psi_{4})\in\Lambda^{2}_{0},$
with $\Lambda^{2}_{0}$ as defined in loc. cit., we conclude that the elements in (3.7), multiplied by $\alpha\in I^{0}_{3}$, are all zero in $\operatorname{H}^{*}(\hat{M})$. By the above arguments, the third equality in (3.6) holds.
Now, we rearrange the above basis of $\operatorname{H}^{*}(\hat{M})$ using the perverse degree instead of the cohomological degree. Note that the generators $\alpha,\beta,\psi_{i}$ all have perverse degree $2$ by [dCHM12, Theorem 4.2.2]. Hence, we obtain
@@
$\operatorname{H}_{\leq 2}^{*}(\hat{M})=\operatorname{H}_{\leq 3}^{*}(\hat{M})=\langle\operatorname{H}_{\leq 0}^{*}(\hat{M}),\alpha,\beta,\psi_{i}\rangle_{1\leq i\leq 4},$
$\operatorname{H}_{\leq 4}^{*}(\hat{M})=\operatorname{H}_{\leq 5}^{*}(\hat{M})=\langle\operatorname{H}_{\leq 2}^{*}(\hat{M}),\alpha^{2},\alpha\beta,\alpha\psi_{i},\psi_{i}\psi_{j}\rangle_{1\leq i<j\leq 4},$
$\operatorname{H}_{\leq 6}^{*}(\hat{M})=\operatorname{H}_{\leq 7}^{*}(\hat{M})=\langle\operatorname{H}_{\leq 4}^{*}(\hat{M}),\alpha^{3},\alpha^{2}\psi_{i},\alpha\gamma\rangle_{1\leq i\leq 4},$
$\operatorname{H}_{\leq 8}^{*}(\hat{M})=\langle\operatorname{H}_{\leq 6}^{*}(\hat{M}),\alpha^{4}\rangle.$
Here we have followed the notation and normalisation conventions of [dCHM12], regarding the perverse filtration on $\operatorname{H}^{*}(\hat{M})$. Precisely:
@@
where $a=\dim(\tilde{B}_{L})=4g(C)-1$.
Combined with the isomorphism $\operatorname{H}^{*}(M)\cong\operatorname{H}^{*}(\hat{M})\otimes\operatorname{H}^{*}(\operatorname{Pic}^{0}(C))$, we obtain
$\sum_{i}\chi({}^{p}\mathcal{H}^{i}({\mathbf{R}h_{*}\operatorname{IC}_{M}}))q^{i}=-(q^{-4}-2q^{-2}+4-2q^{2}+q^{4})(q^{-1/2}+q^{1/2})^{4},$
from which we can read off the integers $n_{g,2[C]}(\operatorname{Tot}_{C}(L))$. The minus sign in the right hand side comes from the fact that $\dim M=8g(C)-3$ is odd. ∎
@@ 4. Higher genus
For simplicity, we write $M=M_{L}(2,1)$ and $\hat{M}=\hat{M}_{L}(2,1)$. By [CL16] (see also [MS23, Theorem 0.4]), we have
${}^{p}\mathcal{H}^{i}(\mathbf{R}h_{*}\operatorname{IC}_{M})\cong\operatorname{IC}(\wedge^{i}\mathbf{R}^{1}\pi_{*}\operatorname{IC}_{\mathcal{C}}),\quad i\in\mathbb{Z},$
where $U\subset\widetilde{B}_{L}$ is the dense open subset parametrising smooth spectral curves and $\pi\colon\mathcal{C}\to U$ is the universal curve. Since a smooth spectral curve has genus $24g(C)-$ 2, the degree of the Laurent polynomial $\sum_{i\in\mathbb{Z}}\chi({}^{p}\mathcal{H}^{i}(\mathbf{R}h_{*}\operatorname{IC}_{M}))q^{i}$ is less than or equal to $24g(C)-$ 2. The coefficient of $q^{4g(C)-2}$ is
@@
To determine the GV invariants for small $g$, recall that we have an isomorphism
@@
From this, we can see that $n_{g,2[C]}(\operatorname{Tot}_{C}(L))=0$ for $g\leq g(C)-1$. Furthermore, we have $n_{g,2[C]}(\operatorname{Tot}_{C}(L))=-\chi(\hat{M})$. This can be computed as follows: Let $P_{t}(M)$ denote the Poincare polynomial of $M$. We can compute $P_{-t}(M)$ by substituting $u=v=-t$ in [CDP11, Equation (A.7)]: